# Adaptive Local Neighborhood-based Neural Networks for MR Image Reconstruction from Undersampled Data

### v2

@@ IV-D Generalizability of LONDN-MRI
Figure 14: Average PSNR on test set (from fastMRI) for LONDN-MRI (MoDL network with UNet denoiser) at 4x k-space undersampling for various overall dataset sizes. Subsets of the dataset are chosen as neighbors in LONDN-MRI at test time. The average PSNR values with DIP, LORAKI, and RAKI, which require no training data are shown as horizontal lines.
@@
PSNR = $\infty$ dB PSNR = 31.45 dB PSNR = 32.15 dB PSNR = 32.72 dB PSNR = 33.15 dB PSNR = 33.22 dB
@@
PSNR = $\infty$ dB PSNR = 33.21 dB PSNR = 36.12 dB PSNR = 36.31 dB PSNR = 36.54 dB PSNR = 36.71
@@
trained with a fixed mask
trained with rand. masks
(2 iterations)
LONDN
@@ VI Conclusions
This paper examined supervised learning of deep unrolled networks at reconstruction time for MRI by exploiting training sets along with local modeling and clustering. We showed advantages for this approach at different k-space undersampling factors over networks learned in a global manner on larger data sets. The training may be connected to a bilevel optimization problem. We also compared different distance metrics for finding neighbors in our approach and regularization to reduce local overfitting. We intend to expand our studies in the future by incorporating non-Cartesian undersampling patterns, such as radial and spiral patterns, as well as deploying them to 3D settings and other imaging modalities. Additionally, the method’s generalizability will be further examined, with a particular emphasis on heterogeneous datasets. We showed benefits for both randomly seeded training of simple models and for fine tuning of sophisticated pre-trained models, and believe our methodology could be applied to a variety of deep learning-based tasks (even beyond image reconstruction) effectively to improve overall performance. Finally, metric learning [47] to improve local clustering and subsequent network adaptation will be an important future direction.

### v3

@@ IV-D Generalizability of LONDN-MRI
Figure 14: Average PSNR on test set (from fastMRI) for LONDN-MRI (MoDL network with UNet denoiser) at 4x k-space undersampling for various dataset sizes. Subsets of the dataset are chosen as neighbors in LONDN-MRI at test time. The average PSNR values with DIP, LORAKI, and RAKI, which require no training data are shown as horizontal lines.
@@
PSNR = $\infty$ dB PSNR = 31.45 dB PSNR = 32.15 dB PSNR = 32.72 dB PSNR = 33.15 dB PSNR = 33.26 dB
@@
PSNR = $\infty$ dB PSNR = 33.21 dB PSNR = 36.12 dB PSNR = 36.31 dB PSNR = 36.54 dB PSNR = 36.77
@@
trained with a fixed mask trained with rand. masks (2 iterations) LONDN
@@ VI Conclusions
This paper examined supervised learning of deep unrolled networks at reconstruction time for MRI by exploiting training sets along with local modeling and clustering. We showed advantages for this approach at different k-space undersampling factors over networks learned in a global manner on larger data sets. The training may be connected to a bilevel optimization problem. We also compared different distance metrics for finding neighbors in our approach and regularization to reduce local overfitting. We intend to expand our studies in the future by incorporating non-Cartesian undersampling patterns, such as radial and spiral patterns, as well as deploying them to 3D settings and other imaging modalities. Additionally, the method’s generalizability will be further examined, with a particular emphasis on heterogeneous datasets. To handle more extreme training-test data variations such as unseen anatomies, we plan to explore patch-based neighbors in local learning schemes for future work. We showed benefits for both randomly seeded training of simple models and for fine tuning of sophisticated pre-trained models, and believe our methodology could be applied to a variety of deep learning-based tasks (even beyond image reconstruction) effectively to improve overall performance. Finally, metric learning [48] to improve local clustering and subsequent network adaptation will be an important future direction.